OpenStudy (anonymous):

Three cell phone towers are shown at the right. The measure if M is 15 less than the measure of K. The measure of L is 5 greater than the measure of M. Which two towers are closest together? Explain

OpenStudy (anonymous):



OpenStudy (kkutie7):

M=k-15 L=5+M

OpenStudy (anonymous):

How is that work? o.O?

OpenStudy (danjs):

Those ar translations of the given info for the angles

OpenStudy (anonymous):

All i understand is M=-15 L=+5

OpenStudy (danjs):

M is 15 less than K --- M = K - 15 L is 5 greater than M --- L = M + 5 when you se "is" that is like 'equals

OpenStudy (danjs):

you can not solve that yet though, 2 equations and 3 variables

OpenStudy (anonymous):

Oh, so its K and L because M is opposite and the smallest angle. Right?

OpenStudy (anonymous):

I'm trying to solve it right now in my paper

OpenStudy (anonymous):

Can you please tell me how to solve it properly?

OpenStudy (danjs):

you can not solve 2 equations with 3 variables, you need at least 3 equations, one for each variable what else can you say about angles of a triangle? they total to 180 degrees M + L + K = 180 try and solve now for M,L and K

OpenStudy (anonymous):

-15+5+x=180

OpenStudy (anonymous):

Three cell phone towers are shown at the right. The measure of M is 15 less than the measure of K. The measure of L is 5 greater than the measure of M. Which two towers are closest together? Explain

OpenStudy (anonymous):



OpenStudy (danjs):

M = K - 15 L = 5 + M 180 = M + L + K ---------------- solve first equation for K K = M + 15 L = 5 + M 180 = M + L + K --------------- Sub in the values of K and L in terms of M and solve for M 180 = M + (5+M) + (M + 15)

OpenStudy (danjs):

M = 160/3 degrees

OpenStudy (anonymous):

53.33

OpenStudy (danjs):

L = 160/3 + 5 K = 160 / 3 + 15

OpenStudy (anonymous):

L=160/3 + 5 = 58.33

OpenStudy (anonymous):

K=160/3 +15 = 68.33

OpenStudy (danjs):

right, then the sides across opposite of an angle are proportional to the coresponding angle measures

OpenStudy (anonymous):

Ok i got it thanks!!

OpenStudy (danjs):

the actual distances cant be found, but the relative distances can since the angles are known

OpenStudy (danjs):

welcome